# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v3

@@ C.2 Proof of Theorem 3
C.3 Krohn-Rhodes Theorem
@@ E.1.2 Details on the evaluated tasks
  • Modular Arithmetic w/ Brackets ($|\Sigma|=12$, $Acc_{rand}=1/5$). Same definition as the modular arithmetic without brackets with a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,),(,[PAD]}\}$. In our experiments $m=5$. An example sequence is as follows:
@@
Table 5: Performance comparison of various recurrent models on regular and context-free language tasks. recurrent models on formal language tasks. We report the median $\pm$ median absolute deviation of 3 independent runs with different random seeds. Scores represent scaled accuracy, with 1.0 indicating perfect performance and 0.0 random guessing. The positive impact of allowing negative eigenvalues ($[-1,1]$ range) versus restricting to positive eigenvalues ($[0,1]$ range) is evident across different model architectures.
@@
Transformer 0.003 $\pm$ 0.013 0.018 $\pm$ 0.009 0.025 $\pm$ 0.000
mLSTM 0.018 $\pm$ 0.035 0.027 $\pm$ 0.013 0.034 $\pm$ 0.001
sLSTM 1.000 $\pm$ 0.000 0.124 $\pm$ 0.000 0.153 $\pm$ 0.020
Mamba $[0,1]$ $0$$\pm$ 0.000 0.066 $\pm$ 0.029 0.072 $\pm$ 0.008
Mamba $[-1,1]$ 1.000 $\pm$ 0.000 0.214 $\pm$ 0.027 0.126 $\pm$ 0.010
DeltaNet $[0,1]$ 0.010 $\pm$ 0.005 0.214 $\pm$ 0.056 0.113 $\pm$ 0.009
DeltaNet $[-1,1]$ 0.999 $\pm$ 0.006 0.826 $\pm$ 0.146 0.129 $\pm$ 0.016

### v5

@@ C.3 Proof of Theorem 3
C.4 Krohn-Rhodes Theorem
@@ E.1.2 Details on the evaluated tasks
  • Modular Arithmetic w/ Brackets, ($|\Sigma|=12$, $Acc_{rand}=1/5$). Same definition as the modular arithmetic without brackets with a set of special tokens $\Sigma_{s}=\{\mathtt{+,-,*,=,),(,[PAD]}\}$. In our experiments $m=5$. An example sequence is as follows:
@@
Table 5: Performance comparison of various recurrent models on regular and context-free language tasks. recurrent models on formal language tasks. We report the median $\pm$ median absolute deviation (left table) and best score (right table) of 3 independent runs with different random seeds. Scores represent scaled accuracy, with 1.0 indicating perfect performance and 0.0 random guessing. The positive impact of allowing negative eigenvalues ($[-1,1]$ range) versus restricting to positive eigenvalues ($[0,1]$ range) is evident across different model architectures.
@@
Mod. Arithm.
(w/ brackets,
no mult)
Transformer 0.003 $\pm$ 0.013 0.018 $\pm$ 0.009 0.064 $\pm$ 0.003 0.025 $\pm$ 0.000
mLSTM 0.018 $\pm$ 0.035 0.027 $\pm$ 0.013 0.114 $\pm$ 0.000 0.034 $\pm$ 0.001
sLSTM 1.000 $\pm$ 0.000 0.124 $\pm$ 0.000 0.163 $\pm$ 0.015 0.153 $\pm$ 0.020
Mamba $[0,1]$ $0$$\pm$ 0.000 0.066 $\pm$ 0.029 0.116 $\pm$ 0.007 0.072 $\pm$ 0.008
Mamba $[-1,1]$ 1.000 $\pm$ 0.000 0.214 $\pm$ 0.027 0.098 $\pm$ 0.009 0.126 $\pm$ 0.010
DeltaNet $[0,1]$ 0.010 $\pm$ 0.005 0.214 $\pm$ 0.056 0.162 $\pm$ 0.018 0.113 $\pm$ 0.009
DeltaNet $[-1,1]$ 0.999 $\pm$ 0.006 0.826 $\pm$ 0.146 0.227 $\pm$ 0.011 0.129 $\pm$ 0.016